Calculate the mole fraction of 5.65g Calcium Chloride with molar mass of 111g/mole 

placed in 15.5g of H2O with molar mass of 18.0g/mole.


1
Expert's answer
2021-03-22T06:13:45-0400

n(CaCl2) = (5.65 g)/(111 g/mol) = 0.051 mol

n(H2O) = (15.5 g)/(18 g/mol) = 0.861 mol

x(CaCl2) = 0.051/(0.051 + 0.861) = 0.0559 = 5.59%